OpenStudy (anonymous):

how do i solve for c in the equation 2sinc(1-cosc) = 4/pi ?

OpenStudy (jdoe0001):

hmmm can you post a picture of the material? I'm assuming that might be a typo =)

OpenStudy (anonymous):

i typed it correctly but i'll post a pic

OpenStudy (anonymous):

\[2[\sin(c)] - \sin(2c) = 4/\pi\]

OpenStudy (anonymous):

then i changed sin(2c) using the double angle formula

OpenStudy (jdoe0001):

I see

OpenStudy (anonymous):

\[2\sin(c) - 2\sin(c)\cos(c)\] = 4/pi

OpenStudy (anonymous):

then i factored out 2sinc

OpenStudy (anonymous):

now i do not know how to solve for c

OpenStudy (anonymous):

if you factored and divided by the constant you may want to take the reciprocal. that way you at least have pi/something instead of something over pi. may be a bit more comprehendable.

OpenStudy (anonymous):

take the reciprocal of what?

OpenStudy (anonymous):

\[2\sin(c) \times (1-cosc) = 4/\pi\]

OpenStudy (anonymous):

\[sinc \times (1-cosc) = 2/\pi\]

OpenStudy (anonymous):

nvm... you can sub sin^2 + cos^2 for 1

OpenStudy (anonymous):

wait a sec...

OpenStudy (anonymous):

i'll wait

OpenStudy (anonymous):

drawing a blank... what class is this for?

OpenStudy (anonymous):

calculus. a blank?

OpenStudy (anonymous):

is this an area or something? what was the original question/setting?

OpenStudy (anonymous):

or a derivative? (by area i meant integral)

OpenStudy (anonymous):

drawing a blank = no good ideas

OpenStudy (anonymous):

so what was the original question that this came from?